# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v2

@@ 7 Conclusion
The asymmetric (weighted) aggregation of prediction sets is a flexible generalization of standard symmetric aggregation. Intuitively, coverage in this setting should vary with the distribution of weights. The results of Vovk and Wang [18] allow us to formalize this intuition for conformal prediction, so that for data-independent weights based on expert priors, we obtain improved guarantees when the weights are sufficiently asymmetric. We extend this framework to data-dependent weights (e.g. weights learned from data), enabling adaptive coverage that reflects the observed input. Experiments on WS coverage and demographic subgroups confirm the practical benefits of this extension, showing that weighted aggregation yields more reliable coverage in challenging settings.
@@ A.1 Proof of Proposition 5
Following the notation of Vovk and Wang [18], we define the merging function
@@
Recall that we use p-variables to represent our p-value functions applied to data; that is, $P_{k}=\widehat{p}_{k}(X,Y)$. Then $\bar{P}=M_{1,\mathbf{v}}(P_{1},\dots,P_{K})$. By Proposition 9 of Vovk and Wang [18], $A_{\mathbf{v}}M_{1,\mathbf{v}}$ is a precise merging function, where $A_{\mathbf{v}}=\min\left\{\frac{1}{v},2\right\}$ and $v=\max\{v_{1},...,v_{K}\}$. Thus, $A_{\mathbf{v}}\bar{P}$ is a p-variable, and $\mathbb{P}\left\{A_{\mathbf{v}}\bar{P}\leq\alpha\right\}\leq\alpha$, or
@@ A.2 Proof of Proposition 5
also leads to coverage of at least $1-\alpha$ for the scaled set, and at least $1-(m^{\prime}+\alpha)$ for the unscaled set. We focus on the correction factor $m^{*}$ because it aligns with the framework of Vovk and Wang [18] and works well in practice.
@@ B.2.2 Local validity over demographic groups
In this section, we evaluate our methods on the Communities and Crimes dataset [37], where the task is to predict the per capita violent crime rate of a community based on its demographic features, and our primary interest is in understanding how coverage differs across communities with varying racial compositions [38].
@@
The demographic groups in Figures 8 and 9 represent communities where a particular racial demographic is in the top $p$-percentile of representation [38]. Figure 8 shows results for $p=50$, and Figure 9 for $p=70$.
@@
Appendix C Other methods for weighted prediction set aggregation
C.1 Weighted majority vote
@@
The idea to combine conformal prediction sets by weighted majority vote was introduced by Gasparin and Ramdas [17] as an extension of the majority vote method first proposed by Cherubin [23]. At first consideration, weighted majority vote appears to differ from our method: weighted majority vote performs weighted aggregation of prediction sets, while our method performs weighted aggregation of the p-values associated with prediction sets. Despite this distinction, Appendix B of Gasparin and Ramdas [17] observes that these two methods are, in fact, dual to each other under data-independent weights, the setting considered in their work. Nevertheless, our p-value formulation enables two key extensions that go beyond weighted majority vote.
@@
  • Our formulation allows us to apply the result of Vovk and Wang [18] to strengthen coverage guarantees for data-independent weights when they are sufficiently asymmetric.
@@
Because our method is a dual formulation to the weighted majority vote method of Gasparin and Ramdas [17], we do not include it as a separate baseline to avoid redundancy.
C.2 Extending the p-variable transformation of Stutz et al. [30] to weighted aggregation
To the best of our knowledge, Gasparin and Ramdas [17] present the only existing method to address weighted prediction set aggregation, and the method of Stutz et al. [30] is designed for the different problem of uncertainty in the ground truth labels. To address their problem, Stutz et al. [30] propose sampling $m$ labels for each calibration point and using the labels to compute $m$ p-values, then taking the unweighted average of these p-values and applying a transformation to obtain a valid p-variable.
Although Stutz et al. [30] only consider the unweighted average, their transformation is general enough to apply to a weighted average of p-values as well, and can therefore be adapted to our setting.
@@
Both our method and the method of Stutz et al. [30] aim to transform a random variable to a p-variable to maintain coverage guarantees. The difference between the two methods lies in the nature of the transformation. Our method applies a linear transformation that preserves the proportions of the weights; this can be important when the weights reflect meaningful quantities, like the weights learned by the routing network of an MoE model. In contrast, Stutz et al. [30] apply a nonlinear, rank-based transformation by computing the empirical CDF of the random variable and returning its value at the observed point. That is, given a random variable $X$, their method estimates its CDF $F$ and uses $F(X)$ as the resulting p-value. While this guarantees validity and preserves ordering, it does not preserve the relative scale between values and therefore discards some of the semantics of the original weights.
@@
To adapt the method of Stutz et al. [30] to this context, we substitute their transformation in place of our linear scaling. We now restate their original setting and transformation in more detail to clarify how their method can be extended to our setting.
The transformation proposed by Stutz et al. [30] was originally developed to address the problem of uncertainty in the ground truth labels. In their setting, each calibration point consists only of an input $X_{i}$; they sample $m$ labels to get $m$ p-values for each calibration point, and then take the unweighted average of these p-values. We denote this unweighted average as $P_{\text{avg}}^{i}$, with distribution function $F$. Their method transforms $P_{\text{avg}}^{i}$ into a valid p-value via $F(P_{\text{avg}}^{i})$, and the prediction set is then the set of all labels such that the corrected p-value $F(P_{\text{avg}}^{i})$ corresponding to each sample is greater than some threshold $\alpha$.
Stutz et al. [30] note that the coverage guarantee using the true CDF holds only with an asymptotic number of samples, as the empirical CDF $\widehat{F}$ approaches the true CDF $F$. To establish a finite-sample guarantee, they introduce a DKW-derived correction $\epsilon$ and define their prediction sets based on $\widehat{F}+\epsilon$. We find that, although their proposed prediction set yields an elegant $(1-\alpha)(1-\delta)$ finite-sample guarantee, the $\epsilon$ correction is extremely conservative in practice.
@@
Figure 11 recreates the Communities and Crimes experiment of Figure 8 with ECDF and ECDF-DKW. Again, we see that ECDF performs very similarly to WA precise on our demographic-conditioned view of Communities and Crimes, with similar coverage (top) and prediction set size (bottom) across all demographics—with ECDF having slightly lower coverage on most demographics, including demographics where WA precise undercovers. Like before, we also observe that ECDF-DKW is so conservative that it has 100% coverage and unbounded prediction sets.
@@
For ECDF-DKW, the finite-sample variant of ECDF, Stutz et al. [30] use DKW to add a finite-sample correction $\epsilon$ to the empirical CDF $\widehat{F}(P_{\text{all}})$, then compare this sum to the significance level $\alpha$. The prediction set with finite-sample guarantees is therefore the set of all labels such that $\widehat{F}(P_{\text{avg}})+\epsilon>\alpha$. However, if $\epsilon$ is already greater than $\alpha$, then this condition is always satisfied and the prediction set includes all labels, becoming unbounded.
@@
Figure 12: Finite-sample correction $\epsilon$ used in ECDF-DKW as a function of the merging set size $|S_{\text{merge}}|$ and user-specified significance level $\delta$. This correction is the offset required to ensure the finite sample guarantee of $(1-\alpha)(1-\delta)$ in Stutz et al. [30]. Dashed lines mark $\alpha$ levels of 0.05 and 0.1. When $\epsilon>\alpha$, the prediction set includes all labels and becomes unbounded, yielding 100% coverage.

### v3

@@ 7 Conclusion
The asymmetric (weighted) aggregation of prediction sets is a flexible generalization of standard symmetric aggregation. Intuitively, coverage in this setting should vary with the distribution of weights. The results of vovk2020combining allow us to formalize this intuition for conformal prediction, so that for data-independent weights based on expert priors, we obtain improved guarantees when the weights are sufficiently asymmetric. We extend this framework to data-dependent weights (e.g. weights learned from data), enabling adaptive coverage that reflects the observed input. Experiments on WS coverage and demographic subgroups confirm the practical benefits of this extension, showing that weighted aggregation yields more reliable coverage in challenging settings.
@@ A.1 Proof of Proposition 5
Following the notation of vovk2020combining, we define the merging function
@@
Recall that we use p-variables to represent our p-value functions applied to data; that is, $P_{k}=\widehat{p}_{k}(X,Y)$. Then $\bar{P}=M_{1,\mathbf{v}}(P_{1},\dots,P_{K})$. By Proposition 9 of vovk2020combining, $A_{\mathbf{v}}M_{1,\mathbf{v}}$ is a precise merging function, where $A_{\mathbf{v}}=\min\left\{\frac{1}{v},2\right\}$ and $v=\max\{v_{1},...,v_{K}\}$. Thus, $A_{\mathbf{v}}\bar{P}$ is a p-variable, and $\mathbb{P}\left\{A_{\mathbf{v}}\bar{P}\leq\alpha\right\}\leq\alpha$, or
@@ A.2 Proof of Proposition 5
also leads to coverage of at least $1-\alpha$ for the scaled set, and at least $1-(m^{\prime}+\alpha)$ for the unscaled set. We focus on the correction factor $m^{*}$ because it aligns with the framework of vovk2020combining and works well in practice.
@@ B.2.2 Local validity over demographic groups
In this section, we evaluate our methods on the Communities and Crimes dataset [redmond2002data], where the task is to predict the per capita violent crime rate of a community based on its demographic features, and our primary interest is in understanding how coverage differs across communities with varying racial compositions [gibbs2023conformal].
@@
The demographic groups in Figures 8 and 9 represent communities where a particular racial demographic is in the top $p$-percentile of representation [gibbs2023conformal]. Figure 8 shows results for $p=50$, and Figure 9 for $p=70$.
@@
B.2.3 Comparing to conservative CQR
In our main results (Figure 3), weighted aggregation typically has higher marginal coverage than CQR (that is, split conformal with CQR nonconformity scores), but also achieves WS coverage that is closer to target. This raises a natural question: does weighted aggregation achieve better WS coverage simply because it is more conservative, or does this reflect genuine adaptivity to difficult regions of the data?
Experimental details
To disentangle these effects, we compare weighted aggregation to a conservative version of CQR. For this baseline, we tune the significance level $\alpha$ of CQR so that its empirical marginal coverage matches that of weighted aggregation as closely as possible. (Note that because conformal prediction with a finite calibration set induces a discrete grid of attainable coverage levels, this matching cannot be exact in general; we select the $\alpha$ that yields the closest available coverage.)
We evaluate both methods over 200 trials and summarize the results in Figure 10. To facilitate comparisons even when marginal coverages are not perfectly matched, we also report $\Delta$ coverage, or the difference between the marginal and WS coverage per trial. [^3] Although potentially unintuitive, empirical WS coverage can occasionally exceed empirical marginal coverage because the WS metric minimizes over a restricted family of slabs (those with at least a $\delta$ fraction of test points, and, in practice, a finite random subset). Consequently, the minimizing slab is not necessarily the region driving undercoverage, and the resulting minimum may be higher than the overall average. For this reason, $\Delta$ coverage can be occasionally less than 0.
Figure 10: Comparison between WA targeted $(0,0.1]$ and CQR at matched empirical marginal coverage. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, $\Delta$ coverage, and prediction set size from left to right. After matching marginal coverage, WA targeted $(0,0.1]$ has more uniform coverage (that is, smaller $\Delta$ coverage) and smaller prediction set size on average.
Discussion of Figure 10
While matching the marginal coverage does move the WS coverage of conservative CQR closer to that of weighted aggregation, performance differences remain. Conservative CQR still undercovers on the WS slab for the Airfoil and Superconduct datasets, and its average $\Delta$ coverage remains consistently larger across all datasets. This indicates that, even when conservative CQR meets the coverage requirement on average (marginally), it fails to maintain that validity locally (on the WS). The lower $\Delta$ coverage of weighted aggregation shows a smaller gap between average and worst-case coverage, indicating a more uniform distribution of validity across the input space.
Meeting this higher marginal coverage also forces conservative CQR to inflate its prediction sets more than weighted aggregation does. In other words, when forced to be equally conservative overall, CQR improves WS coverage primarily by making sets globally wider. In contrast, weighted aggregation improves WS coverage by redistributing uncertainty toward the most challenging regions where coverage failures are more likely. This targeted behavior suggests that weighted aggregation is more adaptive than CQR, rather than simply more conservative.
Appendix C Comparison to other aggregation methods
C.1 Benefits of weighted p-value aggregation
Weighted p-value aggregation is particularly effective in the realistic and practically important scenario when the predictors to be aggregated are not perfectly aligned: that is, when the different predictors capture complementary aspects of a problem, so that no single predictor is uniformly superior (more predictive) across the input space. This setting is represented in Figure 1, and can occur, for example, when the predictors are trained on different modalities or specialize in different operating conditions. In this scenario, an effective aggregation method should (1) faithfully reflect whether the predictors agree or disagree, and (2) adjust to a test point by assigning more weight to the most relevant predictors at that point.
In the standard MoE formulation, weights are optimized for predictive accuracy (under mean squared error or cross-entropy loss), and reflect how relevant each expert is for a given test point. Under this setting, weighted p-value aggregation achieves both of these goals. We elaborate on this below.
C.1.1 p-value aggregation best reflects expert disagreement
One way to taxonomize aggregation methods is to consider whether they combine calibrated uncertainty representations (like p-values or prediction sets) or pre-calibrated representations (like scores). To illustrate how these differ, consider the extreme case of expert diversity where two experts have disjoint prediction sets.
Pre-calibration aggregation methods eliminate disagreement
Methods that combine scores before calibration [luoconformity] convert the experts’ diverse behaviors into a single blended score that no longer encodes which expert supported which region. In the disjoint expert example, averaging pre-calibration scores would collapse the final prediction set into a single region that neither expert truly supports.
To see this more explicitly, let each expert use a standard convex conformal score (e.g. absolute residual or quantile regression score) so each $s_{k}(x,y)$ is convex in $y$. Then, the average score $s_{\text{avg}}(x,y)=\sum_{k}w_{k}s_{k}(x,y)$ is convex, and the conformal set produced from this score is the sublevel set $C_{\text{avg}}=\{y:s_{\text{avg}}(x,y)\leq Q_{\alpha}^{+}\}$, and is also convex. Thus, even if the experts’ individual prediction sets are disjoint, the averaged score produces one connected prediction set which may not reflect the support of either expert.
Set selection methods eliminate disagreement
Methods that select one prediction set [yang2024selection] necessarily discard other experts’ predictions. In the disjoint expert example, the aggregate set reflects only one expert’s prediction, and all complementary evidence from the other expert is lost. This means the aggregate set ignores valid, calibrated support for labels that remain plausible.
p-value aggregation preserves the degree of support from each expert
In the end, the only methods that preserve expert disagreement are the methods that combine, and do not just select, calibrated representations—p-value aggregation, majority vote [cherubin2019majority], and set union. However, these methods differ substantially in how much expert information they retain.
Majority vote and set union operate on binary membership. For weighted majority vote [gasparin2024merging], the prediction set is
$C_{\text{WMV}}(x)=\left\{y:\sum_{k}v_{k}\mathds{1}\{p_{k}(x,y)>\alpha\}\geq\frac{1}{2}\right\},$
so each expert contributes only the indicator $\mathds{1}\{p_{k}(x,y)>\alpha\}$. This tracks whether an expert supports a label, but not how strongly they support it. In contrast, p-value aggregation combines the full p-value functions, with prediction set
$C_{\text{p-value}}(x)=\left\{y:\sum_{k}w_{k}p_{k}(x,y)>\alpha\right\}$
(restated from (5)). If the $i$th expert assigns $p_{i}(x,y)=0.95$ and the $j$th expert assigns $p_{j}(x,y)=0.12$ at $\alpha=0.1$, majority vote records only that both experts include $(x,y)$, while p-value aggregation keeps their relative strength with $w_{i}\cdot 0.95+w_{j}\cdot 0.12$. Thus, p-value aggregation preserves graded expert information, and can distinguish regions of strong support from regions of weak support. This sensitivity to the degree of support from each expert allows p-value aggregation to preserve disagreement more faithfully than binary set-based methods.
Why preserving disagreement matters
With complementary experts, disagreement often reflects genuine multi-modality or heterogeneity in the underlying data. Aggregation methods that collapse this structure risk including labels no expert supports, or excluding labels at least one expert favors. By retaining the full shape and strength of each expert’s p-value function, p-value aggregation ensures that the final set aligns more closely with the true uncertainty represented across experts.
C.1.2 p-value aggregation allows for data-dependent weights
As we see in (1), p-variables satisfy a direct and explicit validity condition. This condition is expressed entirely in terms of the distribution function of a single random variable, and as such, is manipulable: we can combine p-values, examine how the combination behaves relative to this inequality, and apply a correction that restores validity. Prediction sets and nonconformity scores do not admit a similarly explicit validity criterion that allows the aggregated object to be repaired directly after reweighting, so no analogous simple correction exists for majority vote, set union, or score-based averaging.
Consider again the case where experts specialize in different regions of the input space, so that the expert that is most relevant at each point varies across the domain. Majority vote and set union cannot express this form of adaptivity. Even in weighted majority vote, the weights must be fixed in advance or otherwise independent of the test input, so the weighting rule is applied uniformly across the input space. This means every expert contributes its vote at every test point in the same way, regardless of whether that expert is informative for the local region. When an expert is poorly suited to the region around the test point, its vote can (at best) widen the prediction set unnecessarily, or even (at worst) counteract the experts that are actually reliable there.
Set selection methods have a more fundamental limitation. By construction, set selection yields a single global choice of expert, and this remains fixed for all future test points. This is incompatible with any notion of input adaptivity, which is important to settings where local conditions change.
C.2 Comparison to additional adaptive baselines
Additional adaptive baselines
One of the benefits of our proposed method is that its extension to data-dependent weights allows for the weight-aggregated set to adapt to the specific test input. This property of adapting to the test input is called local, adaptive, or conditional.
In our experiments in §6, we compare to conformal quantile regression (CQR) [romano2019conformalized] as our adaptive baseline. CQR is simple but powerful, and can be considered a seminal adaptive method with multiple follow-up works [sesia2020comparison, liu2022conformalized, jensen2022ensemble, rossellini2024integrating]. Our experiments in §6 both compare to CQR (represented as split conformal with CQR scores), and combine our method with CQR (represented as weighted aggregation with CQR scores). To further contextualize the local validity of our approach, we now compare to two additional baselines.
  • Localized conformal prediction (LCP) [guan2019conformal]: LCP reweights calibration scores with weights $w_{i}\propto H(X_{i},X_{n+1})$, where the localization kernel $H$ measures the proximity between the test point $X_{n+1}$ and each calibration point $X_{i}$. At test time, LCP forms an empirical distribution of weighted calibration scores and solves for the significance level $\tilde{\alpha}(x_{n+1})$ that restores marginal coverage after the pointwise weighting. LCP can be computationally intensive, as each test point requires kernel weights over all calibration points.
  • Locally weighted conformal prediction (LWCP) [lei2018distribution]: LWCP normalizes calibration residuals by fitting one model for the mean, a second model for the local error scale, and then rescaling each calibration residual by its estimated variability. Running conformal prediction on these normalized scores yields intervals whose widths adapt to local uncertainty.
We focus on comparing to methods that improve coverage across the input space in a general data-adaptive way. For this reason, we do not compare to methods whose local validity requires the user to first specify a target family of subpopulations or distributional shifts, like gibbs2023conformal.
Weighted aggregation can be combined with other adaptive methods
LCP introduces adaptivity by reweighting calibration scores, and CQR and LWCP introduce adaptivity through adaptive score functions. In contrast, our method introduces adaptivity at the aggregation step, by reweighting the post-calibration p-value functions in a data-dependent way.
When these adaptive methods operate on different steps of the conformal procedure, they generally do not interfere with one another. Thus, because weighted aggregation introduces adaptivity at a different step than LCP, CQR, or LWCP, it complements rather than competes with these methods, and can in principle be combined with any of them.
Figure 11: Comparison between weighted aggregation and adaptive baselines. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, $\Delta$ coverage, and prediction set size from left to right. Although the weighted aggregation methods (WA) are conservative marginally, it also has the most uniform coverage (that is, the smallest $\Delta$ coverage overall), and achieves closest-to-nominal coverage on the WS slab.
Experimental details
For this experiment, we run each of the methods separately to isolate their individual effects. These results are therefore presented differently than in Figure 3, where we had to visualize composition between the absolute residual score / CQR score and split conformal / weighted aggregation. In Figure 11, LCP, CQR, and LWCP are separate methods that are each compared to the weighted aggregation methods (WA) directly. (More precisely, LCP, CQR, and LWCP are each implemented on top of split conformal with the full MoE predictor—each method replacing only the component it introduces, with the rest of the procedure matching split conformal using the full MoE.)
We evaluate all methods over 200 trials and report distributions of marginal coverage, WS coverage, $\Delta$ coverage, and prediction set size (as in Figure 10). Figure 11 summarizes the results.
Discussion of Figure 11
The comparisons between weighted aggregation and the adaptive baselines reflect a trade-off between marginal efficiency and adaptivity. Weighted aggregation is marginally overconservative, whereas the other methods have marginal coverage close to target. However, the weighted aggregation methods consistently brings WS coverage close to target, while the other methods substantially undercover the WS slab. In addition, weighted aggregation almost always achieves the best $\Delta$ coverage: even though its marginal coverage is conservative, the gap between marginal and WS is almost always smaller than the other baselines, indicating that weighted aggregation reallocates uncertainty more effectively in the hardest regions (the WS slab). These results suggest that the main advantages of weighted aggregation are in its adaptivity, by providing more uniform coverage overall and achieving better coverage on the WS slab. We therefore suggest to use weighted aggregation when performance on challenging/minority slices is a priority.
C.3 Weighted majority vote
@@
The idea to combine conformal prediction sets by weighted majority vote was introduced by gasparin2024merging as an extension of the majority vote method first proposed by cherubin2019majority. At first consideration, weighted majority vote appears to differ from our method: weighted majority vote performs weighted aggregation of prediction sets, while our method performs weighted aggregation of the p-values associated with prediction sets. Despite this distinction, Appendix B of gasparin2024merging observes that these two methods are, in fact, dual to each other under data-independent weights, the setting considered in their work. Nevertheless, our p-value formulation enables two key extensions that go beyond weighted majority vote.
@@
  • Our formulation allows us to apply the result of vovk2020combining to strengthen coverage guarantees for data-independent weights when they are sufficiently asymmetric.
@@
Because our method is a dual formulation to the weighted majority vote method of gasparin2024merging, we do not include it as a separate baseline to avoid redundancy.
C.4 Extending the p-variable transformation of stutz2023conformal to weighted aggregation
To the best of our knowledge, weighted majority vote [gasparin2024merging, gasparin2024conformal] is the only existing method to address weighted prediction set aggregation, and the method of stutz2023conformal is designed for the different problem of uncertainty in the ground truth labels. To address their problem, stutz2023conformal propose sampling $m$ labels for each calibration point and using the labels to compute $m$ p-values, then taking the unweighted average of these p-values and applying a transformation to obtain a valid p-variable.
Although stutz2023conformal only consider the unweighted average, their transformation is general enough to apply to a weighted average of p-values as well, and can therefore be adapted to our setting.
@@ C.4 Extending the p-variable transformation of stutz2023conformal to weighted aggregation
Both our method and the method of stutz2023conformal aim to transform a random variable to a p-variable to maintain coverage guarantees. The difference between the two methods lies in the nature of the transformation. Our method applies a linear transformation that preserves the proportions of the weights; this can be important when the weights reflect meaningful quantities, like the weights learned by the routing network of an MoE model. In contrast, stutz2023conformal apply a nonlinear, rank-based transformation by computing the empirical CDF of the random variable and returning its value at the observed point. That is, given a random variable $X$, their method estimates its CDF $F$ and uses $F(X)$ as the resulting p-value. While this guarantees validity and preserves ordering, it does not preserve the relative scale between values and therefore discards some of the semantics of the original weights.
@@
To adapt the method of stutz2023conformal to this context, we substitute their transformation in place of our linear scaling. We now restate their original setting and transformation in more detail to clarify how their method can be extended to our setting.
The transformation proposed by stutz2023conformal was originally developed to address the problem of uncertainty in the ground truth labels. In their setting, each calibration point consists only of an input $X_{i}$; they sample $m$ labels to get $m$ p-values for each calibration point, and then take the unweighted average of these p-values. We denote this unweighted average as $P_{\text{avg}}^{i}$, with distribution function $F$. Their method transforms $P_{\text{avg}}^{i}$ into a valid p-value via $F(P_{\text{avg}}^{i})$, and the prediction set is then the set of all labels such that the corrected p-value $F(P_{\text{avg}}^{i})$ corresponding to each sample is greater than some threshold $\alpha$.
The coverage guarantee for this corrected p-value holds only with an asymptotic number of samples, as the empirical CDF $\widehat{F}$ approaches the true CDF $F$. To establish a finite-sample guarantee, stutz2023conformal introduce a DKW-derived correction $\epsilon$ and define their prediction sets based on $\widehat{F}+\epsilon$. We find that, although their proposed prediction set yields an elegant $(1-\alpha)(1-\delta)$ finite-sample guarantee, the $\epsilon$ correction is extremely conservative in practice.
@@
Figure 13 recreates the Communities and Crimes experiment of Figure 8 with ECDF and ECDF-DKW. Again, we see that ECDF performs very similarly to WA precise on our demographic-conditioned view of Communities and Crimes, with similar coverage (top) and prediction set size (bottom) across all demographics—with ECDF having slightly lower coverage on most demographics, including demographics where WA precise undercovers. As before, we observe that ECDF-DKW is so conservative that it has 100% coverage and unbounded prediction sets.
Figure 14: Finite-sample correction $\epsilon$ used in ECDF-DKW as a function of the merging set size $|S_{\text{merge}}|$ and user-specified significance level $\delta$. This correction is the offset required to ensure the finite sample guarantee of $(1-\alpha)(1-\delta)$ in stutz2023conformal. Dashed lines mark $\alpha$ levels of 0.05 and 0.1. When $\epsilon>\alpha$, the prediction set includes all labels and becomes unbounded, yielding 100% coverage.
@@
For ECDF-DKW, the finite-sample variant of ECDF, stutz2023conformal use DKW to add a finite-sample correction $\epsilon$ to the empirical CDF $\widehat{F}(P_{\text{all}})$, then compare this sum to the significance level $\alpha$. The prediction set with finite-sample guarantees is therefore the set of all labels such that $\widehat{F}(P_{\text{avg}})+\epsilon>\alpha$. However, if $\epsilon$ is already greater than $\alpha$, then this condition is always satisfied and the prediction set includes all labels, becoming unbounded.